OpenStudy (anonymous):

Find all the zeros by factoring each function f(x)=x^3-2x^2+x

OpenStudy (anonymous):

@Hero @jdoe0001 @zepdrix @Nnesha

OpenStudy (anonymous):

\[f(x)=x^3-2x^2+x\]\[0=x^3-2x^2+x\]\[0=x(x^2-2x+1)\]\[0=x(x-1)^2\]

OpenStudy (anonymous):

you can do the rest.

OpenStudy (anonymous):

how did you get the second step?

OpenStudy (anonymous):

It is a perfect square trinomial, (or whatever it is called)

OpenStudy (anonymous):

x^2-2x+1 = x^2 - x - x + 1 = x(x - 1) - (x - 1) = (x-1)(x-1) = (x-1)^2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anytime for you

Nnesha (nnesha):

ohhhh you already got help :(